OpenStudy (anonymous):

what is the square root of x-16=3 solve for x. WHERE x is a real number

OpenStudy (amistre64):

you need to define the question better

OpenStudy (anonymous):

u mean \[\sqrt{x -16}=3 ?\]

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

sqrt(9) = 3 soo... x-16 = 9 x = 25

OpenStudy (mathmagician):

]if you mean \[\sqrt{x}-16=3\] then x=361 else if \[\sqrt{x-16}=3\] then x=25

OpenStudy (anonymous):

THANK YOU SO MUCHHHHH